OpenStudy (anonymous):

How would I go about solving this problem? What is the unit rate for each situation in terms of y to x? Remember to include the appropriate units. y = 2/3x where y is the amount of wall painted and x is the number of hours.

OpenStudy (alexandervonhumboldt2):

y=2/3x x=3/2y. 

OpenStudy (anonymous):

If you multiply both sides by 3, you are left with 3y=2x. So, every two hours, 3 walls are painted.

OpenStudy (anonymous):

Thanks both of you!!